What is meant by saying that a substance is “elastic”? A vertical brass rod of circular section is loaded placing a 5 kg weight on top of it if its length is 50 cm, its radius of cross-section 1 cm, and the Young’s modulus of the material 3.5 x 10^10 Nm-2. Find (a) the contraction of the rod (b) the energy stored in it.
1
Expert's answer
2020-11-02T09:26:02-0500

Solution

When we remove external force on a body or material then it come to it's initial position this type material are called elastic.

a) contraction in rod is given by

"\Delta l=\frac{mgl}{\pi r^2Y}"

Putting value of all variable and constants according to data of question

"\Delta l= \frac{5\times9.8\times50\times10^{-2}}{3.14\times1\times10^{-4}\times3.5\times10^{10}}"

"\Delta l=22.29\times10^{-7}m"

b) stored energy is given

"U=\frac{mg\Delta l}{2}"

"U=\frac{5\times9.8\times22.29\times10^{-7} }{2}"

"U=5.46\times10^{-5} J"